1. Calculate the compound amount when S1500 is deposited in an account earning an annual interest rate of 5%, compounded monthly, for 18 months. 2, How much money should be invested in an account that earns 6% annual interest, com- pounded monthly, in order to have $15, 000 in 5 years? 3. How much interest is earned on a $2000 deposit for 2 years at a 0.12% monthly interest rate. compounded daily?
Answer #1
Principal 1500
time 18 months
interest 5%
Compounding Monthly
Amount= 1500*(1+5%/12)^18
Amount=     1,616.57
Principal P
time 5 Years or 60 months
interest 6%
Compounding Monthly
Amount 15000
Principal= 15000/(1+6%/12)^60
Principal= 11,120.58
Principal 2000
time 2 Years or 24 months
interest 0.120% monthly
Amount= 2000*(1+0.12%)^24
Amount=     2,058.40
Interest          58.40